OpenStudy (anonymous):

slope=5 y-intercept=-2 1/2 write in slope intercept form

OpenStudy (anonymous):

y=mx+b now plug them in

OpenStudy (anonymous):

so its y=5x-5/2 right?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

okay thanks can you please help me on another question please

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

write the equation of the line which passes through the origin and contains (-3,-5)

OpenStudy (anonymous):

did they give you the slop?

OpenStudy (anonymous):

no they did not

OpenStudy (anonymous):

so it passes through the origin which is (0,0) and through the points(-3,-5). 1. first find the slop using y2-y1/x2-x1 2. after that use this formula to find the equation y-y1=m(x-x1)

OpenStudy (anonymous):

i got y=5/3x+0 is that correct??

OpenStudy (anonymous):

i got that too

OpenStudy (anonymous):

oh okay thankyou so so much

OpenStudy (anonymous):

np